from the top of an 80 ft building the angle of elevation of the top of the taller building is 49 degrees and the angle of depression of the base of this building is 62 determine the height of the taller building to the nearest foot


1
Expert's answer
2021-06-24T17:52:19-0400

"let\ d\ be\ the\ distance\ between\ the\ buildings"

"let\ l\ be\ the\ difference\ between\ the\ height\ of\ the\ buildings"

"\tan{\left(62^{\circ}\right)}=\frac{80}{d}=>d=\frac{80}{\tan{\left(62^{\circ}\right)}}"

"\tan{\left(49^{\circ}\right)}=\frac{l}{d}=\frac{l\tan{\left(62^{\circ}\right)}}{80}"

"l=\tan{\left(49^{\circ}\right)}\ast\frac{80}{\tan{\left(62^{\circ}\right)}}=48.9ft"

"height\ of\ the\ taller\ building\ \approx\ 49+80=129 ft"